HETAV PATEL

Grade 11,

Let A be 3x3 matrix such that A'A=I and |A|=I, find the value of |A-I| *(' means transpose) PLZ REPLY SOON...................

Let A be 3x3 matrix such that A'A=I and |A|=I, find the value of |A-I|


*(' means transpose)


PLZ REPLY SOON...................

Grade:12th pass

1 Answers

Sher Mohammad IIT Delhi
10 years ago
See
A^2=I => A^2-I=0
(A-I)(A+I)=0
|A-I||A+I|=0
|A-I|=0

thanks and regards,
Sher Mohammad
B.Tech, IIT Delhi